# Generic flatness of the cohomology of thickenings

### v1

@@ Abstract.
Given a set of $m$ distinct points in projective space over a field, and $t$ a positive integer, a classical question asks for the least degree of a hypersurface that passes through each point with multiplicity at least $t$. Related to this, it remains unresolved whether there exists a dense open set of $m$-tuples of points for which this least degree is constant for each $t\geq 1$. We formulate a conjecture regarding the generic freeness of certain local cohomology modules that would answer this in the affirmative; we then prove this conjecture for up to $n+2$ points in projective $n$-space over a field. In a related direction, we prove a generic flatness result for the cohomology of thickenings of a closed subscheme of projective space over an integral domain.
@@
Y.C.-R. was supported by NSF grant DMS 2502321, and A.K.S. by NSF grants DMS 2101671 and DMS 2349623.
@@ 1. Introduction
(a) A foundational result of Hartshorne [Har66] shows that if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is smooth, where $\mathbb{k}$ has characteristic zero, then the cohomology of the symmetric powers of the conormal bundle exhibits strong vanishing properties; as a consequence, for each $i<\dim X$, the natural map
@@
is an isomorphism for all $t\gg 0$; see [Har66, Theorem 8.1 (b)] and [BBL+19, Theorem 2.8].
(b) This is extended in [BBL+19, Theorem 1.1] as follows: if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is locally a complete intersection, where $\mathbb{k}$ is a field of characteristic zero, then, for each $i<\operatorname{codim}(\operatorname{Sing}X)$, the map () 1.0.1 is an isomorphism for all $t\gg 0$. In particular, if $X$ has at most isolated singular points, then () 1.0.1 is an isomorphism for all $i<\dim X$ and $t\gg 0$.
(c) A version of the Kodaira vanishing theorem for thickenings of locally complete intersection subvarieties of $\mathbb{P}_{\mathbb{k}}^{n}$, for $\mathbb{k}$ of characteristic zero, may be found in [BBL+19, Theorem 1.4]; an asymptotic version of the same holds in positive prime characteristic, [BBL+21, Theorem 1.2].
@@
as $t$ varies: generic flatness guarantees that for a fixed $t$, the cohomology is flat over a dense open subset of $\operatorname{Spec}A$; however, it is not clear whether there is a single dense open subset that suffices for all $t\geq 1$. Our main result provides an affirmative answer to this when $X$ is smooth over $\operatorname{Spec}A$:
@@
Let $A$ be a Noetherian domain containing a field of characteristic zero. Let $X\subset\mathbb{P}^{n}_{A}$ be a closed subscheme that is smooth over $\operatorname{Spec}A$, and $X_{t}$ its $t$-th thickening, as defined above. Fix an integer $j$, and set $\mathcal{F}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{O}_{\mathbb{P}^{n}_{A}}(j)$. Then there exists a nonzero element $a\in A$ such that
@@
The proof of this relies on understanding the asymptotic behavior of the cohomology of powers of the ideal sheaf. Towards this, we use the notion of relatively ample vector bundles; while the theory of ample vector bundles on schemes over a field is well-established following Hartshorne [Har66], we require a relative version over an arbitrary Noetherian base scheme $S$. In Section 2, we develop the necessary theory of $f$-ample vector bundles and establish a fiberwise criterion for ampleness, where $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ is a projective morphism. Our primary example of an $f$-ample vector bundle is the normal bundle $\mathscr{N}_{X}$ associated with a closed immersion $X\lhook\joinrel\longrightarrow\mathbb{P}^{n}_{S}$, assuming that $X$ is smooth over the base $S$.
@@
[…]
@@
Our study of the generic flatness of cohomology modules is motivated by a classical question:
@@
In other words, we ask whether the invariants $\alpha_{t}(X)$ are all constant for a general set of $m$ points in $\mathbb{P}_{\mathbb{k}}^{n}$. Again, a generic flatness argument yields such a dense open subset for a fixed $t$, but the challenge lies in finding a single dense open subset that works for all $t\geq 1$. We propose a conjecture that would settle the above question in the affirmative:
Conjecture 1.3.
Let $\mathbf{a}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left\{a_{ij}\mid 1\leq i\leq m,\,0\leq j\leq n\right\}$ be indeterminates over a field $\mathbb{k}$. Set $A\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbb{k}[\mathbf{a}]$, and let $R$ be the polynomial ring $A[x_{0},\dots,x_{n}]$. Consider the ideal
$I\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathfrak{A}_{1}\cap\dots\cap\mathfrak{A}_{m}$
of $R$, where $\mathfrak{A}_{i}$ is the ideal of size $2$ minors
$\mathfrak{A}_{i}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}I_{2}\begin{pmatrix}x_{0}&x_{1}&\cdots&x_{n}\\ a_{i0}&a_{i1}&\cdots&a_{in}\end{pmatrix}\quad\text{ for }1\leq i\leq m.$
In other words, $I$ is the defining ideal of a generic set of $m$ points in projective space. Set $\mathscr{R}(I)$ to be the Rees algebra $\bigoplus_{t\geq 0}I^{t}T^{t}\subset R[T]$. Then there exists a nonzero element $a\in A$ such that
is a free $A_{a}$-module for each $i\geq 0$.
In Section 4 we prove a special case of this:
@@
Section 1 holds when $m\leq n+2$; in particular, the answer to Section 1 is affirmative for up to $n+2$ points in $\mathbb{P}_{\mathbb{k}}^{n}$.
@@ 2. Ample vector bundles relative to a base scheme
The notion of ample vector bundles on a scheme of finite type over a field was introduced by Hartshorne [Har66]; his main motivation was precisely to study the cohomology of thickenings as in () 1.0.1. For our purposes, we require a relative version of this notion over an arbitrary Noetherian base scheme. While this extension is possibly well-known to experts, we were unable to locate a reference that treats it in a form suitable for our needs. Accordingly, in this short section, we develop the necessary relative theory.
@@
denote the composition of $f$ and $\pi$. By extending the notion of a relatively ample line bundle, see, e.g., [Sta25, Tag 01VG], we obtain the following definition:
@@
For each $s\in S$, [Har70, Proposition III.2.1] implies that the normal bundle to $X_{s}$ in $\mathbb{P}_{\kappa(s)}^{n}$, i.e., $\mathscr{N}_{X}\otimes_{\mathcal{O}_{S}}\kappa(s)$, is an ample vector bundle on $X_{s}$. Therefore $\mathscr{N}_{X}$ is $f$-ample by Theorem 2.2. ∎
Our main interest here lies in the following extension of a well-known vanishing result of Hartshorne, [Har66, Theorem 8.1]; see also [BBL+19, Theorem 2.8]. A key point in the vanishing result below is the existence of a stabilization index that is uniform across all fibers.
@@
Let $S$ be a Noetherian scheme over the field of rational numbers, and $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ be a smooth morphism of relative dimension $d\geq 0$. Let $\mathcal{F}$ be a coherent sheaf on $X$ that is flat over $S$, and set $\mathcal{F}_{s}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{F}\otimes_{\mathcal{O}_{S}}\kappa(s)$ for $s\in S$. Then there exists an integer $t_{0}>0$ such that
@@
Let $s\in S$. By Serre duality, e.g., [Har77, §III.7], the cohomology group displayed above is dual to the cohomology group
$\operatorname{H}^{d-i}\Big(X_{s},\,\left(\operatorname{Sym}^{t}(\mathscr{I}_{X_{s}}/\mathscr{I}_{X_{s}}^{2})\right)^{\vee}\otimes_{\mathcal{O}_{X_{s}}}\mathcal{F}_{s}^{\vee}\otimes_{X_{s}}\omega_{X_{s}}\Big),$
@@
see, e.g., [Eis95, Proposition A2.7]. Consider the coherent sheaf
@@
see [Har66, Lemma 3.1], [Har77, Exercise III.8.4]. Hence, the above spectral sequence degenerates and yields the isomorphism
@@
The cohomology and base change theorem, [Har77, Theorem III.12.11], [Gro63, §7.7], [Mum08, §5], tells us that: (a) $\varphi^{j}(s)$ is an isomorphism if it is surjective; and (b) if $\varphi^{j}(s)$ is surjective, then $\varphi^{j-1}(s)$ is surjective if and only if $R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)$ is locally free in a neighborhood of $s$.
@@ 3. Generic flatness in the smooth setting
    The function $\operatorname{Spec}A\longrightarrow\mathbb{Z}$ where $\mathfrak{p}\longmapsto\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(M(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{-j-n-1}$ is constant for each $i\geq 0$.
@@
The equivalence (a) $\iff$ (c) follows from classical results on cohomological functors, see for example [Gro63, Théorème 7.6.9] or [Har77, Corollary 12.9]. The equivalence (c) $\iff$ (d) may be obtained by applying the graded local duality theorem [BH98, Theorem 3.6.19] to the module $M(\mathfrak{p})$. ∎
@@
Let $A$ be a Noetherian domain containing a field of characteristic zero, and $I$ an ideal of $R=A[x_{0},\dots,x_{n}]$. Then there is a nonzero element $a\in A$ such that
@@
Let $R=A[x_{0},\dots,x_{n}]$ be a standard graded polynomial ring over a Noetherian domain $A$, and $I\subset R$ a homogeneous ideal. Fix $j\in\mathbb{Z}$ and $t\geq 1$. Assume that:
@@
Fix $\mathfrak{p}\in\operatorname{Spec}A$. By [Har77, Exercise II.4.11] or [Gro61a, Proposition 7.1.7], there exists a discrete valuation ring $(V,\mathfrak{n})$ of $Q$ that dominates $A_{\mathfrak{p}}$, i.e., $A_{\mathfrak{p}}\subset V$ and $\mathfrak{p}A_{\mathfrak{p}}=\mathfrak{n}\cap A_{\mathfrak{p}}$. To simplify notation, set $\kappa(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V/\mathfrak{n}$ and $R_{V}(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R_{V}\otimes_{V}\kappa(\mathfrak{n})\cong R_{V}/\mathfrak{n}R_{V}$. Then $\operatorname{H}_{I}^{i}(R_{V})\cong\operatorname{H}_{I}^{i}(R)\otimes_{A}V$ is flat over $V$, equivalently torsionfree over $V$, for each $i\geq 0$. Therefore the injective map
@@
We are now ready to prove the main result:
@@
Let $I\subset R$ be a homogeneous ideal defining $X$ in $\mathbb{P}_{A}^{n}$. By generic freeness, e.g., [Mat89, §24], there exists a nonzero element $a^{\prime}$ in $A$ such that

### v2

@@ Abstract.
We prove a generic flatness result for the cohomology of thickenings of a projective scheme that is smooth over a Noetherian domain containing a field of characteristic zero. Our study is motivated, in part, by a classical question in algebraic geometry: Given a set of $m$ distinct points in projective space over a field, and $t$ a positive integer, determine the least degree of a hypersurface that passes through each point with multiplicity at least $t$. Related to this, it remains unresolved whether there exists a dense open set of $m$-tuples of points for which this least degree is constant for each $t\geq 1$. Investigating this connection in the case of nine points in projective plane, we construct a local cohomology module that is not generically free; moreover, we show that it has infinitely many associated prime ideals.
@@
Y.C.-R. was supported by NSF grant DMS 2502321, and A.K.S. by NSF grants DMS 2101671 and DMS 2349623; E.B. is a member of GNSAGA of INdAM (Italy)
@@ 1. Introduction
(a) A foundational result of Hartshorne [Hart66] shows that if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is smooth, where $\mathbb{k}$ has characteristic zero, then the cohomology of the symmetric powers of the conormal bundle exhibits strong vanishing properties; as a consequence, for each $i<\dim X$, the natural map
@@
is an isomorphism for all $t\gg 0$; see [Hart66, Theorem 8.1 (b)] and [BBLSZ19, Theorem 2.8].
(b) This is extended in [BBLSZ19, Theorem 1.1] as follows: if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is locally a complete intersection, where $\mathbb{k}$ is a field of characteristic zero, then, for each $i<\operatorname{codim}(\operatorname{Sing}X)$, the map () 1.0.1 is an isomorphism for all $t\gg 0$. In particular, if $X$ has at most isolated singular points, then () 1.0.1 is an isomorphism for all $i<\dim X$ and $t\gg 0$.
(c) A version of the Kodaira vanishing theorem for thickenings of locally complete intersection subvarieties of $\mathbb{P}_{\mathbb{k}}^{n}$, for $\mathbb{k}$ of characteristic zero, may be found in [BBLSZ19, Theorem 1.4]; an asymptotic version of the same holds in positive prime characteristic, [BBLSZ21, Theorem 1.2].
@@
as $t$ varies: generic flatness guarantees that for a fixed $t$, the cohomology is flat over a dense open subset of $\operatorname{Spec}A$; however, it is not clear whether there is a single dense open subset that suffices for all $t\geq 1$. Our first main result provides an affirmative answer to this when $X$ is smooth over $\operatorname{Spec}A$:
@@
Let $A$ be a Noetherian domain containing a field of characteristic zero. Let $X\subset\mathbb{P}^{n}_{A}$ be a closed subscheme that is smooth over $\operatorname{Spec}A$, and $X_{t}$ be its $t$-th thickening, as defined above. Fix an integer $j$, and set $\mathcal{F}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{O}_{\mathbb{P}^{n}_{A}}(j)$. Then there exists a nonzero element $a\in A$ such that
@@
The proof of this relies on understanding the asymptotic behavior of the cohomology of powers of the ideal sheaf. Towards this, we use the notion of relatively ample vector bundles; while the theory of ample vector bundles on schemes over a field is well-established following Hartshorne [Hart66], we require a relative version over an arbitrary Noetherian base scheme $S$. In Section 2, we develop the necessary theory of $f$-ample vector bundles and establish a fiberwise criterion for ampleness, where $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ is a projective morphism. Our primary example of an $f$-ample vector bundle is the normal bundle $\mathscr{N}_{X}$ associated with a closed immersion $X\lhook\joinrel\longrightarrow\mathbb{P}^{n}_{S}$, assuming that $X$ is smooth over the base $S$.
@@
$\lim\limits_{\begin{subarray}{c}\longrightarrow\\ t\end{subarray}}\operatorname{Ext}_{R}^{i}\big(R/I^{t},\,R\big)\ =\ \operatorname{H}_{I}^{i}(R),$
@@
The generic freeness of local cohomology modules of certain Rees algebras connects to a classical question regarding points in projective space:
@@
The question above asks whether the invariants $\alpha_{t}(X)$ are all constant for a general set of $m$ points in $\mathbb{P}_{\mathbb{k}}^{n}$. Again, a generic flatness argument yields such a dense open subset for a fixed $t$, but the challenge lies in finding a single dense open subset that works for all $t\geq 1$. In Section 4, we note that an affirmative answer to Section 1 would follow from the generic freeness of local cohomology modules of a Rees algebra; we prove this generic freeness for the case of up to $n+2$ points in projective $n$-space over a field, see Theorem 4.1 (where the answer to Section 1 is known to be affirmative). However, in Section 5, we show that the hoped-for generic freeness fails in the case of nine points in $\mathbb{P}^{2}$. This, in turn, yields a local cohomology module that has infinitely many associated primes — a construction of independent interest (see Theorem 5.1):
@@
[…]
$\mathfrak{A}_{i}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}I_{2}\begin{pmatrix}x_{0}&x_{1}&x_{2}\\ a_{i0}&a_{i1}&a_{i2}\end{pmatrix}$.
Set $I\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathfrak{A}_{1}\cap\dots\cap\mathfrak{A}_{9}\subset R$ and consider the corresponding Rees algebra $\mathscr{R}(I)=\bigoplus_{t\geq 0}I^{t}T^{t}$. Then the following statements hold:
  1.  (a)
    There is no nonzero element $a\in A$ such that $\operatorname{H}_{(x_{0},x_{1},x_{2})}^{2}\left(\mathscr{R}(I)\right)\otimes_{A}A_{a}$ is a flat $A_{a}$-module.
  2.  (b)
    Each of the following sets of associated primes is infinite:
    $\operatorname{Ass}_{A}\left(\operatorname{H}_{(x_{0},x_{1},x_{2})}^{2}\left(\mathscr{R}(I)\right)\right),\quad\operatorname{Ass}_{R}\left(\operatorname{H}_{(x_{0},x_{1},x_{2})}^{2}\left(\mathscr{R}(I)\right)\right),\text{\qquad and \qquad}\operatorname{Ass}_{\mathscr{R}(I)}\left(\operatorname{H}_{(x_{0},x_{1},x_{2})}^{2}\left(\mathscr{R}(I)\right)\right).$
This result highlights why the symbolic powers of an ideal of points have historically been, and remain, notoriously difficult to analyze. Indeed, it shows that the symbolic powers of an ideal of points will inevitably have a non-uniform behavior (see Section 5).
While the question of the finiteness of the Bass numbers of local cohomology modules was raised by Grothendieck [Gro68, Exposé XIII, Conjecture 1.1] and answered in the negative by Hartshorne [Hart69], the question of the finiteness of the set of associated prime ideals is due to Huneke [Hu92, Problem 4]. Affirmative answers include the case of regular rings of prime characteristic [HS93], regular local and affine rings of characteristic zero [Ly93], unramified regular local rings of mixed characteristic [Ly00], smooth $\mathbb{Z}$-algebras [BBLSZ14], and rings of positive prime characteristic with finite Frobenius representation type [TT08, HN17, DQ20]; other positive results over rings of low dimension may be found in [Mar01]. The answer, however, is negative in general; the counterexamples in [Si00, Je21, JS23] come from integer torsion in the local cohomology of finitely generated algebras over the ring of integers, while those in [Ka02, SS04] are for affine algebras over a field. The example constructed in Section 5 is arguably more geometric in flavor: it arises from torsion points in elliptic curves over the complex numbers.
@@ 2. Ample vector bundles relative to a base scheme
The notion of ample vector bundles on a scheme of finite type over a field was introduced by Hartshorne [Hart66]; his main motivation was precisely to study the cohomology of thickenings as in () 1.0.1. For our purposes, we require a relative version of this notion over an arbitrary Noetherian base scheme. While this extension is possibly well-known to experts, we were unable to locate a reference that treats it in a form suitable for our needs. Accordingly, in this short section, we develop the necessary relative theory.
@@
denote the composition of $f$ and $\pi$. By extending the notion of a relatively ample line bundle, see, for example, [Sta26, Tag 01VG], we obtain the following definition:
@@
For each $s\in S$, [Hart70, Proposition III.2.1] implies that the normal bundle to $X_{s}$ in $\mathbb{P}_{\kappa(s)}^{n}$, i.e., $\mathscr{N}_{X}\otimes_{\mathcal{O}_{S}}\kappa(s)$, is an ample vector bundle on $X_{s}$. Therefore $\mathscr{N}_{X}$ is $f$-ample by Theorem 2.2. ∎
Our main interest here lies in the following extension of a well-known vanishing result of Hartshorne, [Hart66, Theorem 8.1]; see also [BBLSZ19, Theorem 2.8]. A key point in the vanishing result below is the existence of a stabilization index that is uniform across all fibers.
@@
Let $S$ be a Noetherian scheme over the field of rational numbers, and $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ be a smooth morphism of relative dimension $d\geq 0$. Let $\mathcal{F}$ be a vector bundle on $X$, and set $\mathcal{F}_{s}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{F}\otimes_{\mathcal{O}_{S}}\kappa(s)$ for $s\in S$. Then there exists an integer $t_{0}>0$ such that
@@
Let $s\in S$. By Serre duality, e.g., [Hart77, §III.7], the cohomology group displayed above is dual to the cohomology group
$\operatorname{H}^{d-i}\Big(X_{s},\,\left(\operatorname{Sym}^{t}(\mathscr{I}_{X_{s}}/\mathscr{I}_{X_{s}}^{2})\right)^{\vee}\otimes_{\mathcal{O}_{X_{s}}}\mathcal{F}_{s}^{\vee}\otimes_{\mathcal{O}_{X_{s}}}\omega_{X_{s}}\Big),$
@@
see, for example, [Eis95, Proposition A2.7]. Consider the vector bundle
@@
see [Hart66, Lemma 3.1], [Hart77, Exercise III.8.4]. Hence, the above spectral sequence degenerates and yields the isomorphism
@@
The cohomology and base change theorem, [Hart77, Theorem III.12.11], [Gro63, §7.7], [Mum08, §5], tells us that: (a) $\varphi^{j}(s)$ is an isomorphism if it is surjective; and (b) if $\varphi^{j}(s)$ is surjective, then $\varphi^{j-1}(s)$ is surjective if and only if $R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)$ is locally free in a neighborhood of $s$.
@@ 3. Generic flatness in the smooth setting
    The function $\operatorname{Spec}A\longrightarrow\mathbb{Z}$ where
    $\mathfrak{p}\longmapsto\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(M(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{-j-n-1}$
    is constant for each $i\geq 0$.
@@
The equivalence (a) $\iff$ (c) follows from classical results on cohomological functors, see for example [Gro63, Théorème 7.6.9] or [Hart77, Corollary 12.9]. The equivalence (c) $\iff$ (d) may be obtained by applying the graded local duality theorem [BH98, Theorem 3.6.19] to the module $M(\mathfrak{p})$. ∎
@@
Let $A$ be a Noetherian domain containing a field of characteristic zero, and $I$ be an ideal of $R=A[x_{0},\dots,x_{n}]$. Then there is a nonzero element $a\in A$ such that
@@
Let $R=A[x_{0},\dots,x_{n}]$ be a standard graded polynomial ring over a Noetherian domain $A$, and $I\subset R$ be a homogeneous ideal. Fix $j\in\mathbb{Z}$ and $t\geq 1$. Assume that:
@@
Fix $\mathfrak{p}\in\operatorname{Spec}A$. By [Hart77, Exercise II.4.11] or [Gro61a, Proposition 7.1.7], there exists a discrete valuation ring $(V,\mathfrak{n})$ of $Q$ that dominates $A_{\mathfrak{p}}$, i.e., $A_{\mathfrak{p}}\subset V$ and $\mathfrak{p}A_{\mathfrak{p}}=\mathfrak{n}\cap A_{\mathfrak{p}}$. To simplify notation, set $\kappa(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V/\mathfrak{n}$ and $R_{V}(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R_{V}\otimes_{V}\kappa(\mathfrak{n})\cong R_{V}/\mathfrak{n}R_{V}$. Then $\operatorname{H}_{I}^{i}(R_{V})\cong\operatorname{H}_{I}^{i}(R)\otimes_{A}V$ is flat over $V$, equivalently torsionfree over $V$, for each $i\geq 0$. Therefore the injective map
@@
We are now ready to prove the first main result:
@@
Let $I\subset R$ be a homogeneous ideal defining $X$ in $\mathbb{P}_{A}^{n}$. By generic freeness, for example, [Mat89, §24], there exists a nonzero element $a^{\prime}$ in $A$ such that